Use the method of substitution to solve the system of linear equations.
9x+4y= 9/4
Y=4x-1
The solution is?
Please show me work, I want to get this!!
Thanks so much!

Found 2 solutions by sachi, Cromlix:
Answer by sachi(548)

9x+4y= 9/4......1
y=4x-1..........2
substituting value of y from 2 in 1
9x+4(4x-1)=9/4
or 9x+16x-4=9/4
or 25x=9/4+4=25/4
or x=1/4
y=4*1/4-1=0

Answer by Cromlix(4381)

9x + 4y = 9/4
y = 4x - 1
Substitute 4x -1 for 'y'
9x + 4(4x - 1) = 9/4
9x + 16x - 4 = 9/4
25x - 4 - 9/4 = 0
Convert -4 to quarters
25x - 16/4 - 9/4 = 0
25x - 25/4 = 0
Common denominator 4
100x - 25 = 0
100x = 25
x = 25/100
x = 1/4
Substitute x = 1/4 into y = 4x -1 => y = 4(1/4) - 1 => y = 0
{ 1/4, 0 )